Felice bought a duplex apartment at a cost of $215,000. Her mortgage payments on the property are $1,020 per month, $516 of which can be deducted from her income taxes. Her real estate taxes total $1,452 per year, and insurance costs $1,224 per year. She estimates that she will spend $1,548 each year per apartment for maintenance, replacing appliances, and other costs. The tenants will pay for all utilities. What must she charge to make $4,600 in profit each year? (Do not round intermediate calculations. Round your answer to 2 decimal places. Ignore any tax effects.) Rent per apartment
Answer #1

Actual cost to be charged on Mortgage payments per Month = 1020-516 = $504

Calculation Of Rent to make profit $4600 each year

Actual cost to be charged on Mortgage payments per Year = 504*12 = 6048

Real estate taxes = 1452

Insurance costs = 1224

Maintenance of apartments @1548 per apartment (1548*2) = 3096

   Total costs = $11820

Target Profit after total cost = $4600

Monthly rent to be charged for Duplex Apartment = 16420/12 = $1368.33

So Rent Per Apartment to be Charged = $684.17